Three times the sum of a number and seven yields the same result as the twice the same number decreased by four.
Answer by checkley71(8403)

3(N+7)=2N-4
3N+21=2N-4
3N-2N=-4-21
N=-25 ANSWER
PROOF
3(-25+7)=2*-25-4
3*-18=-50-4
-54=-54